# FINITE SUBGROUPS OF MAXIMAL ORDER OF THE CREMONA GROUP OVER THE RATIONALS

### v1

@@ Abstract.
Let $\operatorname{Cr}_{\mathbb{Q}}(2)$ be the Cremona group of rank $2$ over rational numbers. We give a classification of large finite subgroups $G$ of $\operatorname{Cr}_{\mathbb{Q}}(2)$ and give a new sharp bound smaller (but not multiplicative) than $M(\mathbb{Q})=120960=2^{7}\cdot 3^{3}\cdot 5\cdot 7$; the one given in [Ser09]. In particular, we prove that any finite subgroup $G\subset\operatorname{Cr}_{\mathbb{Q}}(2)$ has order $\mid G\mid\leq 432$ and Lemma 4.1 provides a group of order $432$. We use the modern approach of minimal $G-$surfaces, given a (smooth) rational surface $S\subset\mathbb{P}^{2}$ defined over $\mathbb{Q}$, we study the finite subgroups $G\subset\operatorname{Aut}_{\mathbb{Q}}(S)$ of automorphisms of $S$. We give the best bound for the order of $G\subset\operatorname{Aut}(S)$ for surfaces with a conic bundle structure invariant by $G$. We also give the best bound for the order of $G\subset\operatorname{Aut}_{\mathbb{Q}}(S)$ for all rational Del Pezzo surfaces of some given degree. In addition, we give descriptions of the finite subgroups of automorphisms of conic bundles and Del Pezzo surfaces of maximal size.
@@ 1. Introduction
The group $\operatorname{Cr}_{k}(1)$ is the group of automorphisms of the projective line, and hence it is isomorphic to the projective linear group $\operatorname{PGL}(2,k)$. The classification of finite subgroups of $\operatorname{PGL}(2)$ is well-known and goes back to F. Klein. In the case $n=2$ and $k=\mathbb{C}$, the group $\operatorname{Cr}_{\mathbb{C}}(2)$ is the plane Cremona group over the field of complex numbers. The classification of finite subgroups of $\operatorname{Cr}_{\mathbb{C}}(2)$ has been studied and has a long history, going back to the $19$th century (see the references in [Bla06] and [DI09]).
In this paper, we consider the plane Cremona group over the field of rational numbers, denoted by $\operatorname{Cr}_{\mathbb{Q}}(2)$. This group $\operatorname{Cr}_{\mathbb{Q}}(2)$ has been studied in [Ser09]. The main theorem in [Ser09] gives a sharp multiplicative bound for the orders of the finite subgroups $G$ of $\operatorname{Cr}_{\mathbb{Q}}(2)$, which is $M(\mathbb{Q})=120960=2^{7}\cdot 3^{3}\cdot 5\cdot 7$.
Our goal is to give a classification of finite subgroups $G$ of $\operatorname{Cr}_{\mathbb{Q}}(2)$ and give a new sharp bound smaller (not multiplicative) than the one given in [Ser09]. Our main theorem is the following:
@@
As already used in [Ser09], every finite subgroup $G\subset\operatorname{Cr}_{\mathbb{Q}}(2)$ is conjugate, by a birational map $\phi:\mathbb{P}^{2}_{\mathbb{Q}}\dashrightarrow S$, to a finite subgroup of biregular automorphisms $\phi G\phi^{-1}\subset\operatorname{Aut}_{\mathbb{Q}}(S)$. Moreover, one can reduce to the case where S is a Del Pezzo surface or where the group preserves a conic bundle. The degree of a Del Pezzo being between $1$ and $9$ and since every finite subgroup of automorphisms of a Del Pezzo surface of degree $7$ is conjugate to a finite subgroup of $\operatorname{Aut}(\mathbb{P}^{2})$, we get $9$ cases to consider. We give the upper bound for the order of a finite group in each of the cases in the next table, and give references for an example, and the proof that it is maximal.
@@
$G-$surface structure of $G$ order example lemma conic bundles $D_{6}\times D_{6}$ $144$ 3.5 3.4 $dP_{9}$ $(\mathbb{Z}/2\mathbb{Z})^{2}\rtimes\operatorname{Sym}_{3}$ $24$ 6.13 6.12 $dP_{8}$ $(D_{6}\times D_{6})\rtimes\mathbb{Z}/2\mathbb{Z}$ $288$ 9.2 9.3 $dP_{6}$ $(\mathbb{Z}/6\mathbb{Z})^{2}\rtimes D_{6}$ $432$ 4.1 4.2 $dP_{5}$ $\operatorname{Sym}_{5}$ $120$ 5.2 5.3 $dP_{4}$ $(\mathbb{Z}/2\mathbb{Z})^{4}\rtimes\operatorname{Sym}_{3}$ $96$ 7.5 7.6 $dP_{3}$ $\operatorname{Sym}_{5}$ $120$ 8.6 8.5 $dP_{2}$ $(\mathbb{Z}/2\mathbb{Z})^{3}\rtimes\operatorname{Sym}_{3}$ $48$ 10.2 10.1 $dP_{1}$ $D_{6}$ $12$ 11.2 11.1
@@ 2. Preliminaries
We use the modern approach to the problem initiated in the works of Yuri Manin and the second author (see a survey of their results in [Isk96]). Their work gives a clear understanding of the conjugacy problem via the concept of a rational $G-$surface: a pair $(G,S)$ in which $S$ is a (rational smooth projective) surface and $G\subset\operatorname{Aut}(S)$ is a group acting biregularly on $S$.
@@
A birational map $\varphi:S\dashrightarrow\mathbb{P}^{2}$ yields an isomorphism between $\operatorname{Cr}(S)$ and $\operatorname{Cr}(\mathbb{P}^{2})$. The Cremona group is then isomorphic to the group of birational maps of any rational surface. We say that two groups $G\subset\operatorname{Cr}(S)$ and $G^{\prime}\subset\operatorname{Cr}(S^{\prime})$ are birationally conjugate if there exists a birational map $\varphi:S\dashrightarrow S^{\prime}$ such that $\varphi G\varphi^{-1}=G^{\prime}$. This means that the two groups represent the same conjugacy class in the Cremona group and conversely a conjugacy class of a finite subgroup $G$ of $\operatorname{Cr}_{\mathbb{C}}(2)$ can be realized as a birational isomorphism class of $G-$surfaces. In this way classification of conjugacy classes of subgroups of $\operatorname{Cr}_{\mathbb{C}}(2)$ becomes equivalent to the birational classification of $G-$surfaces. On another way, any subgroup of $\operatorname{Bir}(S)$ may therefore be viewed as a subgroup of the Cremona group, up to conjugacy. A $G-$equivariant analog of a minimal surface allows one to concentrate on the study of minimal $G-$surfaces, $i.e.$, surfaces which cannot be $G-$equivariantly birationally and regularly mapped to another $G-$surface. Minimal $G-$surfaces turn out to be $G-$isomorphic either to a conic bundle or a Del Pezzo surface of degree $d$ where $1\leq d\leq 9$.
@@
We will study the first case in Section 3 and the second case in Sections $4-11$.
@@ 3. Rational Surfaces with Conic Bundle structure
Let $(S,\pi)$ and $(\tilde{S},\tilde{\pi})$ be two conic bundles. We say that $\varphi:S\dashrightarrow\tilde{S}$ is a birational map of conic bundles if $\varphi$ is a birational map which sends a general fibre of $\pi$ on a general fibre of $\tilde{\pi}$. This corresponds to ask for $\alpha\in\operatorname{Aut}(\mathbb{P}^{1})$ such that $\tilde{\pi}\circ\varphi=\alpha\circ\pi$.
We say that a conic bundle $(S,\pi)$ is minimal if any birational morphism of conic bundles $(S,\pi)\rightarrow(\tilde{S},\tilde{\pi})$ is an isomorphism.
@@
$(\mathbb{G}_{m})^{2}\times(\mathbb{A}^{2}\setminus\{0\})^{2}$$\rightarrow$ $(\mathbb{A}^{2}\setminus\{0\})^{2},$
@@
The class of $(x_{0},x_{1},y_{0},y_{1})$ will be written $[x_{0}:x_{1};y_{0}:y_{1}]$. The projection
$\tau_{n}:\mathbb{F}_{n}\rightarrow\mathbb{P}^{1},~{}[x_{0}:x_{1};y_{0}:y_{1}]\mapsto[y_{0}:y_{1}]$
identifies $\mathbb{F}_{n}$ with $\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}}(n)\oplus\mathcal{O}_{\mathbb{P}^{1}})$ as a $\mathbb{P}^{1}-$bundle over $\mathbb{P}^{1}$.
The disjoint sections $s_{n},s_{-n}\subset\mathbb{F}_{n}$ given by $x_{0}=0$ and $x_{1}=0$ have self-intersection $n$ and $-n$, respectively. The fibres $f\subset\mathbb{F}_{n}$ given by $y_{0}=0$ and $y_{1}=0$ are linearly equivalent and of self-intersection 00. We moreover get $\mathrm{Pic}(\mathbb{F}_{n})=\mathbb{Z}f\oplus\mathbb{Z}s_{-n}=\mathbb{Z}f\oplus\mathbb{Z}s_{n}$.
@@
We contract orbits of components in any singular fibre, and assume that the action of $g$ to be minimal. By [Bla09, Lemma 3.8], any singular fibre of $\pi$ is twisted by some element of $\langle g\rangle$. If $g$ twists a singular fibre, then ($g$ is an involution) and the set $\mathrm{Fix}(g)$ of fixed locus of $g$ is a $2:1$ cover of $\mathbb{P}^{1}$ via $\pi$ ([Bla09, Lemma $6.1$]). So no curve contained in a fiber is fixed point-wise by $g$. Otherwise, there is no singular fiber and by [Bla09, Lemma 3.2.], the surface $S$ is isomorphic to the Hirzebruch surface $\mathbb{F}_{n}$, for some integer $n\geq 0$. If $n=0$, then the action is given by
$[x_{0}:x_{1};y_{0}:y_{1}]\mapsto[a\cdot x_{0}+b\cdot x_{1}:c\cdot x_{0}+d\cdot x_{1};y_{0}:y_{1}],$
@@
$[x_{0}:x_{1};y_{0}:y_{1}]\mapsto[x_{0}:\lambda\cdot x_{1}+P_{n}(y_{0},y_{1})\cdot x_{0};y_{0}:y_{1}],$
@@
Let $S$ be a rational surface defined over $\mathbb{Q}$ and $\pi:S\rightarrow C$ be a conic bundle. If $G$ is a finite subgroup of $\operatorname{Aut}_{\mathbb{Q}}(S,\pi)$, then $\mid G\mid$ divides $144$.
@@
Let $S$ be a rational surface defined over $\mathbb{Q}$ and $\pi:S\rightarrow C$ be a conic bundle. Let $G$ be a finite subgroup of $\operatorname{Aut}_{\mathbb{Q}}(S,\pi)$. Since the surface $S$ is rational over $\mathbb{Q}$, then it has rational points and $S(\mathbb{Q})$ is Zariski dense. The curve $C\subset\mathbb{P}^{2}$ is a smooth rational curve defined over $\mathbb{Q}$ as $C(\mathbb{Q})$ is not empty and $C\simeq\mathbb{P}^{1}$ over $\mathbb{Q}$. We have a natural homomorphism $\bar{\pi}:G\rightarrow\operatorname{Aut}(\mathbb{P}^{1}_{\mathbb{Q}})=\operatorname{PGL}(2,\mathbb{Q})$ that satisfies $\bar{\pi}(g)\pi=\pi g$, for every $g\in G$. Let $G^{\prime}=\mathrm{ker}\bar{\pi}$, hence, we have the following exact sequence
(1) $1\rightarrow G^{\prime}\rightarrow G\rightarrow\bar{\pi}(G)\subseteq\operatorname{Aut}(\mathbb{P}^{1}_{\mathbb{Q}})\rightarrow 1$
Hence, $\mid\bar{\pi}(G)\mid\subseteq\operatorname{Aut}(\mathbb{P}^{1}_{\mathbb{Q}})=\operatorname{PGL}(2,\mathbb{Q})$ and by [Bea10, Theorem 4.2.], $\mid\bar{\pi}(G)\mid$ divides $12$. By Lemma 3.3, any element $g\in G^{\prime}\setminus\{id\}$ acts non-trivially on any smooth fibre, this implies that $G^{\prime}\subseteq\operatorname{Aut}(\mathbb{P}^{1}_{\mathbb{Q}})=\operatorname{PGL}(2,\mathbb{Q})$. By [Bea10, Theorem 4.2.], $\mid G^{\prime}\mid$ divides $12$. Since $\mid G\mid=\mid G^{\prime}\mid\cdot\mid\bar{\pi}(G)\mid$, then $\mid G\mid$ divides $12\cdot 12=144$. ∎
@@
Let $S=\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\pi:S\rightarrow\mathbb{P}^{1}$ be the projection on a factor, that is a conic bundle. Any subgroup of $\operatorname{Aut}_{\mathbb{Q}}(S,\pi)\subset\operatorname{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ is isomorphic to a subgroup $G$ of $\operatorname{PGL}_{2}(\mathbb{Q})\times\operatorname{PGL}_{2}(\mathbb{Q})$. Let $G\subset\operatorname{PGL}_{2}(\mathbb{Q})$ be a finite subgroup, then the group $H:=G\times G$ is a finite subgroup of $\operatorname{PGL}_{2}(\mathbb{Q})\times\operatorname{PGL}_{2}(\mathbb{Q})$, hence, the subgroup of $\operatorname{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ isomorphic to $H$ preserves the fibration through $\pi$. Moreover, if $G\cong D_{6}$, then $H$ is of order $144$.
4. Del Pezzo surfaces of degree $6$
Let $S$ be a Del Pezzo surface of degree $6$, defined over $\mathbb{Q}$. Let $X$ be the blow-up of $\mathbb{P}^{2}$ at the three points [1:0:0],[0:1:0][1:0:0],~{}[0:1:0][ 1 : 0 : 0 ] , [ 0 : 1 : 0 ] and $[0:0:1]$, therefore, $X$ is the standard Del Pezzo surface given in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ given by

### v2

@@ Abstract.
Let $\operatorname{Cr}_{\mathbb{Q}}(2)$ be the Cremona group of rank $2$ over rational numbers. we give a classification of large finite subgroups $G$ of $\operatorname{Cr}_{\mathbb{Q}}(2)$ and give a new sharp bound smaller (but not multiplicative) than $M(\mathbb{Q})=120960=2^{7}\cdot 3^{3}\cdot 5\cdot 7$; the one given in [22]. In particular, we prove that any finite subgroup $G\subset\operatorname{Cr}_{\mathbb{Q}}(2)$ has order $\mid G\mid\leq 432$ and Lemma 4.1 provides a group of order $432$. We use the modern approach of minimal $G$-surfaces, given a (smooth) rational surface $S\subset\mathbb{P}^{2}$ defined over $\mathbb{Q}$, we study the finite subgroups $G\subset\operatorname{Aut}_{\mathbb{Q}}(S)$ of automorphisms of $S$. We give the best bound for the order of $G\subset\operatorname{Aut}(S)$ for surfaces with a conic bundle structure invariant by $G$. We also give the best bound for the order of $G\subset\operatorname{Aut}_{\mathbb{Q}}(S)$ for all rational Del Pezzo surfaces of some given degree. In addition, we give descriptions of the finite subgroups of automorphisms of conic bundles and Del Pezzo surfaces of maximal size.
@@ 1. Introduction
The group $\operatorname{Cr}_{k}(1)$ is the group of automorphisms of the projective line, and hence it is isomorphic to the projective linear group $\operatorname{PGL}(2,k)$. The classification of finite subgroups of $\operatorname{PGL}(2)$ is well-known and goes back to F. Klein. In the case $n=2$ and $k=\mathbb{C}$, the group $\operatorname{Cr}_{\mathbb{C}}(2)$ is the plane Cremona group over the field of complex numbers. The classification of finite subgroups of $\operatorname{Cr}_{\mathbb{C}}(2)$ has been studied and has a long history, going back to the $19$th century (see the references in [4] and [8]).
In this paper, we consider the plane Cremona group over the field of rational numbers, denoted by $\operatorname{Cr}_{\mathbb{Q}}(2)$. This group $\operatorname{Cr}_{\mathbb{Q}}(2)$ has been studied in [22]. The main theorem in [22] gives a sharp multiplicative bound for the orders of the finite subgroups $G$ of $\operatorname{Cr}_{\mathbb{Q}}(2)$, which is $M(\mathbb{Q})=120960=2^{7}\cdot 3^{3}\cdot 5\cdot 7$, this means that the order of any finite subgroup of $\operatorname{Cr}_{\mathbb{Q}}(2)$ divides $2^{7}\cdot 3^{3}\cdot 5\cdot 7$.
Our goal is to give a classification of finite subgroups $G$ of $\operatorname{Cr}_{\mathbb{Q}}(2)$ and give a new sharp bound smaller (not multiplicative) than the one given in [22]. Our main theorem is the following:
@@
As already used in [22], every finite subgroup $G\subset\operatorname{Cr}_{\mathbb{Q}}(2)$ is conjugate, by a birational map $\phi:\mathbb{P}^{2}_{\mathbb{Q}}\dashrightarrow S$, to a finite subgroup of biregular automorphisms $\phi G\phi^{-1}\subset\operatorname{Aut}_{\mathbb{Q}}(S)$. Moreover, one can reduce to the case where S is a Del Pezzo surface or where the group preserves a conic bundle. The degree of a Del Pezzo being between $1$ and $9$ and since every finite subgroup of automorphisms of a Del Pezzo surface of degree $7$ is conjugate to a finite subgroup of $\operatorname{Aut}(\mathbb{P}^{2})$, we get $9$ cases to consider. We give the best upper bound for the order of a finite group, an example, and the proof that it is maximal in each of the cases.
@@
$G$-surface structure of $G$ order example lemma
conic bundles $D_{6}\times D_{6}$ $144$ 3.6 3.5
$dP_{9}$ $(\mathbb{Z}/2\mathbb{Z})^{2}\rtimes\operatorname{Sym}_{3}$ $24$ 6.13 6.12
$dP_{8}$ $(D_{6}\times D_{6})\rtimes\mathbb{Z}/2\mathbb{Z}$ $288$ 9.2 9.3
$dP_{6}$ $(\mathbb{Z}/6\mathbb{Z})^{2}\rtimes D_{6}$ $432$ 4.1 4.2
$dP_{5}$ $\operatorname{Sym}_{5}$ $120$ 5.2 5.3
$dP_{4}$ $(\mathbb{Z}/2\mathbb{Z})^{4}\rtimes\operatorname{Sym}_{3}$ $96$ 7.5 7.6
$dP_{3}$ $\operatorname{Sym}_{5}$ $120$ 8.6 8.5
$dP_{2}$ $(\mathbb{Z}/2\mathbb{Z})^{3}\rtimes\operatorname{Sym}_{3}$ $48$ 10.2 10.1
$dP_{1}$ $D_{6}$ $12$ 11.2 11.1
@@ 2. Preliminaries
We use the modern approach to the problem initiated in the works of Yuri Manin and the second author (see a survey of their results in [13]). Their work gives a clear understanding of the conjugacy problem via the concept of a rational $G$-surface.
We consider groups acting on a surface. A $G$-surface, denote by $(G,S)$, is a pair in which $S$ is a (rational) surface and $G\subset\operatorname{Aut}(S)$ is a group acting biregularly on $S$.
@@
A birational map $\varphi:S\dashrightarrow\mathbb{P}^{2}$ yields an isomorphism between $\operatorname{Cr}(S)$ and $\operatorname{Cr}(\mathbb{P}^{2})$. The Cremona group is then isomorphic to the group of birational maps of any rational surface. We say that two groups $G\subset\operatorname{Cr}(S)$ and $G^{\prime}\subset\operatorname{Cr}(S^{\prime})$ are birationally conjugate if there exists a birational map $\varphi:S\dashrightarrow S^{\prime}$ such that $\varphi G\varphi^{-1}=G^{\prime}$. The choice of $\varphi$ does not change the conjugacy class of the group and any element of the conjugacy class can be obtained in this manner. The converse is true, if $G$ is finite ((see for example a proof in [6, Theorem $1.4$])). This means that the two groups represent the same conjugacy class in the Cremona group and conversely a conjugacy class of a finite subgroup $G$ of $\operatorname{Cr}_{\mathbb{C}}(2)$ can be realized as a birational isomorphism class of $G$-surfaces. In this way classification of conjugacy classes of subgroups of $\operatorname{Cr}_{\mathbb{C}}(2)$ becomes equivalent to the birational classification of $G$-surfaces. On another way, any subgroup of $\operatorname{Bir}(S)$ may therefore be viewed as a subgroup of the Cremona group, up to conjugacy. A $G$-equivariant analog of a minimal surface allows one to concentrate on the study of minimal $G$-surfaces, $i.e.$, surfaces which cannot be $G$-equivariantly birationally and regularly mapped to another $G$-surface. Minimal $G$-surfaces turn out to be $G$-isomorphic either to a conic bundle or a Del Pezzo surface of degree $d$ where $1\leq d\leq 9$.
@@
We will study the first case in Section $3$ and the second case in Sections 4114-11.
@@ 3. Rational Surfaces with Conic Bundle structure
[…]
@@
$(\mathbb{G}_{m})^{2}\times(\mathbb{A}^{2}\setminus\{0\})^{2}$ $\longrightarrow$ $(\mathbb{A}^{2}\setminus\{0\})^{2}$
@@
The class of $(x_{0},x_{1},y_{0},y_{1})$ will be written $x_{0}:x_{1};y_{0}:y_{1}]$. The projection
$\tau_{n}:\mathbb{F}_{n}\longrightarrow{\mathbb{P}^{1}},~[x_{0}:x_{1};y_{0}:y_{1}]\mapsto[y_{0}:y_{1}]$
identifies $\mathbb{F}_{n}$ with $\mathbb{P}_{{\mathbb{P}^{1}}}(\mathcal{O}_{{\mathbb{P}^{1}}}\oplus\mathcal{O}_{{\mathbb{P}^{1}}}(n)$ as a $\mathbb{P}^{1}$}-bundle over $\mathbb{P}^{1}$}.
The disjoint sections $s_{n},s_{-n}\subset\mathbb{F}_{n}$ given by $x_{0}=0$ and $x_{1}=0$ have self-intersection $n$ and n-n, respectively. The fibres $f\subset\mathbb{F}_{n}$ given by $y_{0}=0$ and $y_{1}=0$ are linearly equivalent and of self-intersection 0. We moreover get ${\mathrm{Pic}}(\mathbb{F}_{n})=\mathbb{Z}f\oplus\mathbb{Z}s_{-n}=\mathbb{Z}f\oplus\mathbb{Z}s_{n}$.
We have the following result:
Proposition 3.3.
Minimal rational surfaces (([2, Theorem V.10]))
Let $S$ be a minimal rational surface. Then $S$ is isomorphic to $\mathbb{P}^{2}$ or to one of the surfaces $\mathbb{F}_{n}$, with $n\neq 1$.
@@
We contract orbits of components in any singular fibre, and assume that the action of $g$ to be minimal. By [5, Lemma 3.8], any singular fibre of $\pi$ is twisted by some element of $\langle g\rangle$. If $g$ twists a singular fibre, then ($g$ is an involution) and the set $\mathrm{Fix}(g)$ of fixed locus of $g$ is a $2:1$ cover of $\mathbb{P}^{1}$ via $\pi$ ([5, Lemma $6.1$]). So no curve contained in a fiber is fixed point-wise by $g$. Otherwise, there is no singular fiber and by [5, Lemma 3.2.], the surface $S$ is isomorphic to the Hirzebruch surface $\mathbb{F}_{n}$, for some integer $n\geq 0$. If $n=0$, then the action is given by
$x_{0}:x_{1};y_{0}:y_{1}]\mapsto[a\cdot x_{0}+b\cdot x_{1}:c\cdot x_{0}+d\cdot x_{1};y_{0}:y_{1}],$
@@
$x_{0}:x_{1};y_{0}:y_{1}]\mapsto[x_{0}:\lambda\cdot x_{1}+P_{n}(y_{0},y_{1})\cdot x_{0};y_{0}:y_{1}],$
@@
Let $S$ be a rational surface defined over $\mathbb{Q}$ and $\pi:S\rightarrow\mathbb{P}^{1}$ be a conic bundle. If $G$ is a finite subgroup of $\operatorname{Aut}_{\mathbb{Q}}(S,\pi)$, then $\mid G\mid$ divides $144$.
@@
Let $S$ be a rational surface defined over $\mathbb{Q}$ and $\pi:S\rightarrow\mathbb{P}^{1}$ be a conic bundle. Let $G$ be a finite subgroup of $\operatorname{Aut}_{\mathbb{Q}}(S,\pi)$. We have a natural homomorphism $\bar{\pi}:G\rightarrow\operatorname{Aut}(\mathbb{P}^{1}_{\mathbb{Q}})=\operatorname{PGL}(2,\mathbb{Q})$ that satisfies $\bar{\pi}(g)\pi=\pi g$, for every $g\in G$. Let $G^{\prime}=\mathrm{ker}\bar{\pi}$, hence, we have the following exact sequence
(1) $1\rightarrow G^{\prime}\rightarrow G\rightarrow\bar{\pi}(G)\subseteq\operatorname{Aut}(\mathbb{P}^{1}_{\mathbb{Q}})\rightarrow 1.$
Hence, $\mid\bar{\pi}(G)\mid\subseteq\operatorname{Aut}(\mathbb{P}^{1}_{\mathbb{Q}})=\operatorname{PGL}(2,\mathbb{Q})$ and by [3, Theorem 4.2.], $\mid\bar{\pi}(G)\mid$ divides $12$. After extending the scalar to $\bar{\mathbb{Q}}$, we apply Lemma 3.4, then any element $g\in G^{\prime}\setminus\{id\}$ acts non-trivially on any smooth fibre, this implies that $G^{\prime}\subseteq\operatorname{Aut}(\mathbb{P}^{1}_{\mathbb{Q}})=\operatorname{PGL}(2,\mathbb{Q})$. By [3, Theorem 4.2.], $\mid G^{\prime}\mid$ divides $12$. Since $\mid G\mid=\mid G^{\prime}\mid\cdot\mid\bar{\pi}(G)\mid$, then $\mid G\mid$ divides $12\cdot 12=144$. ∎
@@
Let $S=\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\pi:S\rightarrow\mathbb{P}^{1}$ be the projection on a factor, that is a conic bundle. Any subgroup of $\operatorname{Aut}_{\mathbb{Q}}(S,\pi)\subset\operatorname{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ is isomorphic to a subgroup $H$ of $\operatorname{PGL}_{2}(\mathbb{Q})\times\operatorname{PGL}_{2}(\mathbb{Q})$. Let $G\subset\operatorname{PGL}_{2}(\mathbb{Q})$ be a finite subgroup, then the group $H:=G\times G$ is a finite subgroup of $\operatorname{PGL}_{2}(\mathbb{Q})\times\operatorname{PGL}_{2}(\mathbb{Q})$, hence, the subgroup of $\operatorname{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ isomorphic to $H$ preserves the fibration through $\pi$. Moreover, if $G\cong D_{6}$ (a dihedral group of order $12$), then $H$ is of order $144$.
4. Del Pezzo surfaces of degree 66
Let $S$ be a Del Pezzo surface of degree $6$, defined over $\mathbb{Q}$. Let $X$ be the blow-up of $\mathbb{P}^{2}$ at the three points [1:0:0],[0:1:0][1:0:0],~[0:1:0] and $[0:0:1]$, therefore, $X$ is the standard Del Pezzo surface given in $\mathbb{P}^{2}\times\mathbb{P}^{2}$ given by